# Gravitational Wave Signal Denoising and Merger Time Prediction By Deep Neural Network

### v1

@@ V.1 Signal denoising
Fig 2 presents the denoising results for different chirp masses. It can be observed that our model demonstrates considerable denoising performance. The distribution of the results is shown in Fig 3. Since 99% of the denoised data achieve an overlap 0.9 with the template data, we have set the displayed overlap range between 0.9 and 1 for ease of visualization. The results for the dataset with a chirp mass of are slightly inferior to those for other datasets. Specifically, the SNR of some denoised data is significantly lower than that of the corresponding template data. This phenomenon occurs because, for smaller chirp masses, the pre-merger inspiral and approach processes are slower, resulting in less pronounced amplitude variations across the entire data segment. This substantially confuses the model’s judgment.
In Tab 2, we present detailed statistical results. It can be observed that our model precisely recovers the phase for data across different chirp masses. Overall, 91.98% of the data, post-denoising, achieves an overlap 0.99 with the template data, demonstrating our model’s sensitivity to phase recovery. Regarding amplitude recovery, while performance is slightly lower for datasets with small chirp mass, the overall results indicate good amplitude recovery, with 90.52% of the denoised data showing an SNR difference 1 compared to the template data.
Chirp mass | Overlap | SNR differences |
@@ V.2 Signal detection
When dealing with real GW data streams, it is essential to determine when a GW signal is detected in order to obtain subsequent predictions of the merger time. Given the capability of our denoising network to recover GW signals from noise, we have established a workflow, as illustrated in Fig 4, to detect GW signals in GW data streams. The denoising network processes data within a 30-day window, which moves across the GW data stream in hourly steps, outputting the corresponding denoised waveforms. Consequently, there will be overlapping segments between consecutive windows, and we calculate the overlap of these segments. When the overlap exceeds a predefined threshold, it is considered that a GW signal has been detected.
In this test, we will generate 5,000 data pairs for each dataset corresponding to different chirp masses. These pairs consist of current moment data and data from one hour earlier. The SNR for these data pairs is uniformly sampled between 10 and 50. Each dataset contains half GW data and half pure noise data. According to the design of our workflow, we calculated the overlaps for the data pairs in each dataset. To explore the selection of a reasonable threshold, we calculated the true positive rate (TPR) and false positive rate (FPR) for each dataset, as shown in Fig 5. We selected 0.978 as the threshold. At this value, the FPR for all datasets is zero, and the TPR is close to 1. This indicates that our model exhibits high sensitivity to GW signals at this threshold while not responding to pure noise. In Fig 6, we present the confusion matrices for each dataset at the selected threshold. Even for a chirp mass of , only 34 GW signals were missed, resulting in a detection accuracy of 98.64%. Furthermore, the detection accuracy increases with higher chirp masses. These results demonstrate the robustness and rationality of our workflow.
@@ V.3 Merger time prediction
To further explore the overall performance of our model, we conducted a statistical analysis of the results across the entire test set based on the time distance to the merger event. The results, shown in Tab 3, indicate that the average prediction error decreases and the distribution becomes more stable as the dataset approaches closer to the merger time. The time to merger of our dataset ranges from 0 days to 10 days, with 89.1% of the data achieving a prediction error of 24 hours. This demonstrates that our model is capable of making reasonably accurate predictions of merger times well in advance of the event.
@@ VI Concluding remarks
Advance warnings for MBHB mergers are valuable for possible multi-messenger observations with EM and GW detections. However, the overwhelming noise present in the space-based GW detection, combined with the slow accumulation of SNR in the inspiral phase, makes this task particularly challenging. This challenge arises mainly from the need to process long-duration signals, which introduce significant complexity into model training. Short data windows capture detectable features only near the merger phase, significantly limiting early warning capabilities. To address this, we integrate data dimensionality reduction and feature separation techniques to denoise signals up to 30 days long, allowing sufficient SNR to accumulate before the merger. This enalbles efficient detection during the inspiral phase of MBHBs. Our work also has implications for detecting other faint GW signals. Additionally, our model predicts merger times based on denoised waveforms, facilitating more precise planning for multi-messenger observations of MBHB’s EM and GW signals from its merger phase.
We evaluated our model in three key scenarios: signal denoising, signal detection, and merger time prediction. For signal denoising, our model effectively restored the amplitude and phase of noisy signals, achieving an overlap 0.99 with the signal templates for 91.98% of the denoised samples, and an SNR difference 1 for 90.52% of the samples, demonstrating the effectiveness of our denoising network. For signal detection, we designed a workflow to differentiate between pure noise and GW signals. We explored optimal threshold settings to avoid false responses to noise while maintaining high sensitivity to GW signals, achieving a detection accuracy exceeding 98.64% on our test dataset. In terms of merger time prediction, our model demonstrated considerable prediction accuracy; for data occurring no more than 10 days before the merger, the prediction error generally did not exceed 24 hours.
In the future, we aim to enhance the model’s capability to process data with low chirp masses, including improvements in training strategies and architectural design. We also plan to test the robustness of our model confront with possible data anomalies in future works. Space-based GW detection missions may encounter data anomalies and non-stationarities, such as data gaps caused by routine maintenance or hardware issues [83, 60], or acceleration glitches in the GRS system [84]. Moreover, the waveform model used in this study, IMRphenomD, does not include higher-order modes, and we leave these more realistic challenges for future exploration.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China under Grant Nos. 2021YFC2201901, 2021YFC2201903, 2020YFC2200601, 2020YFC2200901, and 2021YFC2203004, as well as the National Science Foundation of China (NSFC) under Grant Nos. 12405076, 12347103, and 12247187. Additionally, this work is also supported by the International Partnership Program of Chinese Academy of Sciences, No. 025GJHZ2023106GC.References
@@
- Abbott et al. [2021] R. Abbott, T. Abbott, S. Abraham, F. Acernese, K. Ackley, A. Adams, C. Adams, R. X. Adhikari, V. Adya, C. Affeldt, et al., Tests of general relativity with binary black holes from the second ligo-virgo gravitational-wave transient catalog, Physical review D 103, 122002 (2021).
@@
- Luo and Mesgarani [2019] Y. Luo and N. Mesgarani, Conv-tasnet: Surpassing ideal time–frequency magnitude masking for speech separation, IEEE/ACM transactions on audio, speech, and language processing 27, 1256 (2019).
- Luo et al. [2020a] Y. Luo, Z. Chen, and T. Yoshioka, Dual-path rnn: efficient long sequence modeling for time-domain single-channel speech separation, in ICASSP 2020-2020 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP) (IEEE, 2020) pp. 46–50.
@@
- Luo et al. [2020b] Z. Luo, Z. Guo, G. Jin, Y. Wu, and W. Hu, A brief analysis to taiji: Science and technology, Results in Physics 16, 102918 (2020b).

### v2

@@ V.1 Signal denoising
Fig 3 presents the denoising results for different chirp masses. It can be observed that our model demonstrates considerable denoising performance. The distribution of the results is shown in Fig 4. Since 99% of the denoised data achieve an overlap 0.9 with the template data, we have set the displayed overlap range between 0.9 and 1 for ease of visualization. The results for the dataset with a chirp mass of are slightly inferior to those for other datasets. This phenomenon occurs because, for smaller chirp masses, the pre-merger inspiral and approach processes are slower, resulting in less pronounced amplitude variations and a less steep frequency variation across the entire data segment. This substantially confuses the model’s judgment.
In Tab 2, we present detailed statistical results. It can be observed that our model precisely recovers the phase for data across different chirp masses. Overall, 91.98% of the data, post-denoising, achieves an overlap 0.99 with the template data, demonstrating our model’s sensitivity to phase recovery. Regarding amplitude recovery, while performance is slightly lower for datasets with small chirp mass, the overall results indicate good amplitude recovery, with 95.82% of the denoised data has an SNR loss 5% compared to the template SNR.
Chirp mass | Overlap | SNR loss (%) |
@@ V.2 Signal detection
In this test, we will generate 5,000 data pairs for each dataset corresponding to different chirp masses. These pairs consist of current moment data and data from one hour earlier. The SNR for these data pairs is uniformly sampled between 10 and 50. Each dataset is a balanced binary classification dataset, with half of the samples being GW signals and the other half being pure noise. According to the design of our workflow, we calculated the overlaps for the data pairs in each dataset. To explore the selection of a reasonable threshold, we calculated the true positive rate (TPR) and false positive rate (FPR) for each dataset, as shown in Fig 5. We selected 0.978 as the threshold. At this value, the FPR for all datasets is zero, and the TPR is close to 1. This indicates that our model exhibits high sensitivity to GW signals at this threshold while not responding to pure noise. In Fig 6, we present the confusion matrices for each dataset at the selected threshold. Even for a chirp mass of , only 34 GW signals were missed, resulting in a detection accuracy of 98.64%. Furthermore, the detection accuracy increases with higher chirp masses. These results demonstrate the robustness and rationality of our workflow.
@@ V.3 Merger time prediction
To further explore the overall performance of our model, we conducted a statistical analysis of the results across the entire test set based on the time distance to the merger event. The results, shown in Tab 3, indicate that the average prediction error decreases and the distribution becomes more stable as the dataset approaches closer to the merger time. The time to merger of our dataset ranges from 0 days to 10 days, with 89.1% of the data achieving a absolute prediction error of 24 hours. This demonstrates that our model is capable of making reasonably accurate predictions of merger times well in advance of the event.
@@ VI Concluding remarks
Advance warnings for MBHB mergers are valuable for possible multi-messenger observations with EM and GW detections. However, the SNR of MBHB signals accumulates in a highly nonlinear fashion, with the final hour before merger contributing as much as 99% of the total SNR [54]. As a result, the inspiral segment exhibits a significantly lower SNR, which makes detection challenging. This challenge arises mainly from the need to process long-duration signals, which introduce significant complexity into model training. Short data windows capture detectable features only near the merger phase, significantly limiting early warning capabilities. To address this, we integrate data dimensionality reduction and feature separation techniques to denoise signals up to 30 days long, allowing sufficient SNR to accumulate before the merger. This enalbles efficient detection during the inspiral phase of MBHBs. Our work also has implications for detecting other faint GW signals. Additionally, our model predicts merger times based on denoised waveforms, facilitating more precise planning for multi-messenger observations of MBHB’s EM and GW signals from its merger phase.
We evaluated our model in three key scenarios: signal denoising, signal detection, and merger time prediction. For signal denoising, our model effectively restored the amplitude and phase of noisy signals, achieving an overlap 0.99 with the signal templates for 91.98% of the denoised samples, and an SNR loss 5% for 95.82% of the samples, demonstrating the effectiveness of our denoising network. For signal detection, we designed a workflow to differentiate between pure noise and GW signals. We explored optimal threshold settings to avoid false responses to noise while maintaining high sensitivity to GW signals, achieving a detection accuracy exceeding 98.64% on our test dataset. In terms of merger time prediction, our model demonstrated considerable prediction accuracy; for data occurring no more than 10 days before the merger, the absolute prediction error generally did not exceed 24 hours.
In the future, we aim to enhance the model’s capability to process data with low chirp masses, including improvements in training strategies and architectural design. Furthermore, in this work, the current method is only applicable to the detection of a single inspiral signal. We are currently conducting research on the event rate of multiple inspiral signal overlaps detected by the Taiji detector. Once we have reached definitive conclusions, we plan to extend the research to the detection of multiple inspiral signal overlaps. We believe this work will serve as a solid starting point. We also plan to test the robustness of our model confront with possible data anomalies in future works. Space-based GW detection missions may encounter data anomalies and non-stationarities, such as data gaps caused by routine maintenance or hardware issues [91, 66], or acceleration glitches in the GRS system [92]. Moreover, the waveform model used in this study, IMRphenomD, does not include higher-order modes, and we leave these more realistic challenges for future exploration.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China under Grant Nos. 2021YFC2201901, 2021YFC2201903, 2020YFC2200601, 2020YFC2200901, and 2021YFC2203004, as well as the National Science Foundation of China (NSFC) under Grant Nos. 12405076, 12347103, and 12247187. Additionally, this work is also supported by the International Partnership Program of Chinese Academy of Sciences, No. 025GJHZ2023106GC. We thank the anonymous referee for valuable comments and suggestions that helped us to improve the manuscript.References
@@
- Abbott et al. [2021a] R. Abbott, T. Abbott, S. Abraham, F. Acernese, K. Ackley, A. Adams, C. Adams, R. X. Adhikari, V. Adya, C. Affeldt, et al., Tests of general relativity with binary black holes from the second ligo-virgo gravitational-wave transient catalog, Physical review D 103, 122002 (2021a).
@@
- Abbott et al. [2021b] R. Abbott, T. Abbott, S. Abraham, F. Acernese, K. Ackley, A. Adams, C. Adams, R. Adhikari, V. Adya, C. Affeldt, et al., Gwtc-2: compact binary coalescences observed by ligo and virgo during the first half of the third observing run, Physical Review X 11, 021053 (2021b).
- Abbott et al. [2023] R. Abbott, T. Abbott, F. Acernese, K. Ackley, C. Adams, N. Adhikari, R. Adhikari, V. Adya, C. Affeldt, D. Agarwal, et al., Gwtc-3: Compact binary coalescences observed by ligo and virgo during the second part of the third observing run, Physical Review X 13, 041039 (2023).
@@
- Dewdney et al. [2009] P. E. Dewdney, P. J. Hall, R. T. Schilizzi, and T. J. L. Lazio, The square kilometre array, Proceedings of the IEEE 97, 1482 (2009).
- Abell et al. [2009] P. A. Abell, J. Allison, S. F. Anderson, J. R. Andrew, J. R. P. Angel, L. Armus, D. Arnett, S. Asztalos, T. S. Axelrod, S. Bailey, et al., Lsst science book, version 2.0, (2009).
- Nandra et al. [2013] K. Nandra, D. Barret, X. Barcons, A. Fabian, J.-W. d. Herder, L. Piro, M. Watson, C. Adami, J. Aird, J. M. Afonso, et al., The hot and energetic universe: A white paper presenting the science theme motivating the athena+ mission, arXiv preprint arXiv:1306.2307 (2013).
- Palenzuela et al. [2010] C. Palenzuela, L. Lehner, and S. L. Liebling, Dual jets from binary black holes, Science 329, 927 (2010).
@@
- Luo et al. [2020a] Z. Luo, Z. Guo, G. Jin, Y. Wu, and W. Hu, A brief analysis to taiji: Science and technology, Results in Physics 16, 102918 (2020a).
@@
- Baghi [2022] Q. Baghi, The lisa data challenges, arXiv preprint arXiv:2204.12142 (2022).
- Kirillov et al. [2023] A. Kirillov, E. Mintun, N. Ravi, H. Mao, C. Rolland, L. Gustafson, T. Xiao, S. Whitehead, A. C. Berg, W.-Y. Lo, et al., Segment anything, in Proceedings of the IEEE/CVF International Conference on Computer Vision (2023) pp. 4015–4026.
- Dosovitskiy [2020] A. Dosovitskiy, An image is worth 16x16 words: Transformers for image recognition at scale, arXiv preprint arXiv:2010.11929 (2020).
- Luo et al. [2020b] Y. Luo, Z. Chen, and T. Yoshioka, Dual-path rnn: efficient long sequence modeling for time-domain single-channel speech separation, in ICASSP 2020-2020 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP) (IEEE, 2020) pp. 46–50.